OpenStudy (anonymous):

How do I solve this system?? Y=-x+3 Y=3/2x-2

OpenStudy (lynfran):

equate the 2 equations and solve..

OpenStudy (anonymous):

1/2x+1?

OpenStudy (lynfran):

now bring the x" to one side of the equation...like.. -x+x+3+2=3/2x+x-2+2 3+2=3/2x+x

OpenStudy (anonymous):

No this doesn't make sense

OpenStudy (anonymous):

All I know is the answer is 2,1 but idk how to get it

OpenStudy (lynfran):

\[-x+3=\frac{ 3 }{ 2 }x-2\]\[-x+3+2=\frac{ 3 }{ 2 }x\]\[-x+5=\frac{ 3 }{ 2 }x\]\[5=\frac{ 5 }{ 2 }x\]\[5/\frac{ 5 }{ 2 }=x\]\[2=x\]

OpenStudy (lynfran):

now sub x=2 into equation 1 or 2 to get y value

OpenStudy (lynfran):

using equation 1 we get y=-x+3 y=-(2)+3 y=-2+3 y=1

OpenStudy (lynfran):

so yes the answer is (2,1)